OpenStudy (anonymous):

The equation 2w+5j=60 is used to determine the number of water bottles w and the number of juice bottles j that can be bought fo $60. I f you purchase 4 bottles of juice, how many bottles of water can you buy?

OpenStudy (anonymous):

Help?

OpenStudy (anonymous):

can you help me person viewing my thing well question

OpenStudy (pjarmy1234):

I am looking in my pre cal book for an example.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so lets see we are solving for w which is water

OpenStudy (anonymous):

oh i see what you do is you - 5 on both sides of the equation

OpenStudy (anonymous):

then you divide 2 on both sides of the equation

OpenStudy (anonymous):

then you have w alone

OpenStudy (pjarmy1234):

Right. I am trying substitution.

OpenStudy (anonymous):

wait that was wrong

OpenStudy (anonymous):

\[ 2w+5j=60\] and you are told \(j=4\) so you get \[2w+20=60\] or \[2w=40\] then divide by 2 to find \(w\)

OpenStudy (anonymous):

ok i am goijg to plug in the answer choices and see what i get which are A. 10 B. 15 C. 20 and D. 25

OpenStudy (anonymous):

its not A

OpenStudy (pjarmy1234):

I got 15

OpenStudy (anonymous):

is it 8 bottles of water? is it correct?

OpenStudy (anonymous):

j = 4 2w + 5j = 60 now sub 4 in for j 2w + 5(4) = 60 2w + 20 = 60 2w = 60 - 20 2w = 40 (divide by 2 to get w by itself) 2w/2 = 40/2 w = 20 if you purchase 4 bottles of juice, then you are purchasing 20 bottles of water, to equal a total of 60.

OpenStudy (anonymous):

sorry :) i got 5w instead of 2w. my bad

OpenStudy (pjarmy1234):

Well, that makes sense. The 20.

OpenStudy (anonymous):

yea it is not B either

OpenStudy (hba):

2w+5j=60 [plug j=4]

OpenStudy (hba):

Then kaboom the value of 'w' will come :)

OpenStudy (anonymous):

or you can plug in the answer choice take alot of work but it is easier

OpenStudy (anonymous):

You need to learn how to do it without answer choices because what if the question doesn't supply answer choices.

OpenStudy (hba):

Well you can go for this when you have answer options and check LHS=RHS if yes then your answer option is correct if not move on to the next :)

OpenStudy (hba):

If you wanna do it and you do not have the options 2w+5j=60 [plug j=4]

OpenStudy (anonymous):

yea that will work to and it is easy and quick

OpenStudy (anonymous):

thank you

OpenStudy (hba):

You are welcome :)